# Methodology for Fine-Grain GPU Power Visibility and Insights

### v1

@@ Abstract
Ubiquity of AI coupled with its steep power demands make optimizing GPU power a priority as large GPU-based clusters are often employed to train and serve AI models. An important first step in optimizing GPU power consumption is high-fidelity and fine-grain power measurement of key AI computations on GPUs. To this end, we observe that as GPUs get more powerful, the resulting sub-millisecond to millisecond executions make fine-grain power analysis challenging. In this work, we first carefully identify the challenges in obtaining fine-grain GPU power profiles. To address these challenges, we devise FinGraV methodology where we employ execution time binning, careful CPU-GPU time synchronization, and power profile differentiation to collect fine-grain GPU power profiles across prominent AI computations and across spectrum of scenarios. Using FinGraV power profiles, we make several observations pertaining to GPU power variation over executions and over time, GPU sub-component power consumptions across different scenarios, and power behavior over interleaved executions of multiple computations. Equipped with these observations, we conclude with several recommendations to optimize the power for these ubiquitous accelerators.
@@ I Introduction
A key enabler for the current AI wave has been the compute and memory horsepower availed by modern GPUs. As demand for AI sky-rockets, hyperscalers continue to deploy large-scale GPU clusters to train and serve AI models [1, 2]. This increasing demand for AI computation translates into increasing AI energy needs, which is dominated by power consumption of GPUs. As such, optimizing GPU power and energy consumption can go a long way in reducing energy consumption and capital costs of large-scale GPU clusters being deployed for AI. GPU power optimization is not only beneficial for AI but also for other domains, as the use cases for which GPUs are being utilized continue to widen. As an example, even for the HPC domain, the majority (more than 75%) of node-level power provisioning for Hewlett Packard Enterprise Frontier [3], the world’s first exascale supercomputer, is for GPUs.
An important first step in optimizing GPU power is better visibility into GPU power. Specifically, in this work, we focus on measuring fine-grain GPU power profiles along both time and space dimensions. That is, first, for fine-grain profiling along time, we focus on power measurements throughout the execution of a GPU computation (commonly referred to as a kernel in GPU parlance). Second, for fine-grain profiling along space dimension, we focus on breaking down GPU power consumption into sub-components (e.g., compute cores, HBM memory, etc.). We believe that such fine-grain visibility can go a long way in identifying avenues for and hence design of intelligent GPU power management capabilities.
@@
However, fine-grain GPU power analysis is challenging for many reasons as depicted in Figure 1. First, as GPUs get more powerful (higher compute/memory throughputs), execution times for computations can often end up in sub-millisecond (ms) to few ms range. This makes getting fine-grain power profiles challenging even with a 1ms power sampler natively available on a GPU (C1) [4, 5, 6]. Further, a common technique of repeated kernel executions is challenging to use as kernels with such low execution times can manifest execution time variation (e.g., due to slight differences in memory allocation and hence access patterns) which must be tackled (C2). Not only do execution time variations need to be accounted for, even power variations can occur across executions, across interleaving of various kinds of computations and more (C3). Finally, using a high frequency GPU-side power sampler can either lead to repeated CPU-GPU communication for online measurements (as CPU schedules computations on the GPU today) or require careful synchronization of CPU-GPU time during post-processing of power logs obtained to identify logs that belong to the kernel execution (C4).
To tackle these challenges and provide fine-grain GPU power visibility, in this work, we propose FinGraV(Fine-Grain Visibility) methodology to collect fine-grain power profiles on the state-of-the-art AMD Instinct MI300X GPU, which is being used exclusively to serve all Llama 405B live traffic at Meta [7]. To design FinGraV, we harness GPU-side power logging and via careful synchronization of CPU-GPU time, we identify power logs of interest along with identifying its specific time in the kernel’s execution, thus getting a fine-grain profile in time. Further, we employ execution time binning strategy to tackle execution time variation. Finally, we study power behavior for a given GPU kernel under different scenarios to identify power variations that can manifest and identify specific power profiles of interest for a specific kernel.
We study the profiles that FinGraV leads to for prominent AI computations (matrix-matrix multiplication and collective communication kernels) across a spectrum of scenarios. We make several observations based on these profiles about GPU power variation over executions and over time, GPU sub-component power consumptions across different scenarios, and power behavior over interleaved executions of multiple operators. Finally, based on these observations, we conclude with a set of recommendations to optimize GPU power via both software and hardware optimizations.
@@
  • To address the above identified challenges, we propose FinGraV (Fine-Grain Visibility) methodology, where we employ execution time binning, careful CPU-GPU time synchronization, and power profile differentiation to collect fine-grain GPU power profiles.
@@
  • From the above FinGraV profiles, we make several observations about GPU power variation over executions and over time, GPU sub-component power consumptions across different scenarios, and power behavior over interleaved executions of multiple computations.
@@ II-A AMD Instinct MI300X GPU and Sub-components
We focus in this work on fine-grain power analysis of the state-of-the-art AMD Instinct MI300X GPU, based on AMD CDNA 3 architecture, shown in Figure 2. As depicted, the MI300X GPU is a chiplet-based design which harnesses advanced packaging to integrate heterogeneous chiplets each specialized for a specific function. Reading the figure from the top, MI300X has accelerator complex dies (XCD) vertically stacked over I/O dies (IOD), which in turn are stacked over a passive silicon interposer. There are a total of four IODs each having two XCDs stacked for a total of eight XCDs in a single MI300X GPU [8, 9].
@@
Figure 3: Challenges in doing fine-grain GPU power analysis.
@@ II-B ML Focus and Operators of Interest
While AI workloads comprise a variety of operators, we consciously focus on two primary operators for our analysis, namely, general matrix-matrix multiplications (GEMM) kernels and communication kernels for they contribute to the majority of the AI execution time [13]. We consider both GEMM and communication kernels across a spectrum of scenarios (GEMMs: compute versus memory-bound; communication: latency versus bandwidth-bound) and setups (isolated, interleaved executions, and more) in order to study their effect on power consumption.
@@ III Challenges to Fine-grain GPU Power Analysis
Low sampling frequency: As compute throughput and memory bandwidth made available by high-end GPUs continue to scale, short of problem sizes scaling commensurately, GPU kernel execution times can often end up in sub-millisecond (ms) to few ms range. This is certainly true for the AI kernels we have benchmarked on MI300X GPU (Section V). This means that any CPU-side GPU power measurements at low sampling frequency, as Figure 3a depicts, can completely miss sampling power for a given kernel.
@@
Power variance across executions: Finally, note that the above execution time variation, coupled with the fact that power is continuous rather than discrete, leads to power variation manifesting as well. Specifically, as depicted in Figure 3d, we observe in this work that repeated executions of the same kernel (tagged 1/2/3 in Figure 3d) or interleaving of a kernel with other kernels (kernel 1 interleaved with kernel x and y) can lead to different power profiles. As such, identifying which power profiles to focus on is important.
@@ IV-A Addressing Challenges
We discuss in this section, the broad strokes of our solution to the challenges we identified in Section III.
@@
On-GPU power logger: We harness a 1ms power logger available internally at AMD on MI300X. We discuss using the proposed FinGraV methodology in tandem with externally available power logging tools on AMD GPUs such as amd-smi [14] in Section VI.
@@
High-resolution CPU-GPU sync: As the internal power logger logs power measurements on the GPU while being agnostic of kernel start/end events as discussed in Section III, we need to employ careful syncing of CPU-GPU time to identify the power log which was taken during the execution of the kernel (LOI) and where in the kernel execution was the log taken (TOI). To do so, first, we read a GPU timestamp counter before kernel execution from the CPU side. Second, we separately benchmark the delay to read this timestamp. Finally, along with power logging, we also log this timestamp value (depicted in Figure 4b). By associating the GPU timestamp with every power log and syncing a GPU timestamp (T0) with a specific timestamp on the CPU (Tc), we can post-process the power logs to identify LOI/TOI by identifying the kernel-start/end times (also in CPU time domain) in relation to the synced CPU time Tc.
@@
Kernel execution time binning: Note that, as discussed in Section III, a single run is insufficient to create fine-grain power profiles. That said, correlating measurements across runs is challenging due to kernel execution time variation. To tackle this, we employ a simple strategy of execution time binning as depicted in Figure 4c. That is, based on our empirical experiments, we guide the user towards two heuristics for fine-grain power profiles: (1) #runs to execute and (2) margin of execution time variation to allow. By excluding outlier runs with considerable variations and fine-tuning the margin to kernel execution time (Section V-B), we lower the effects of execution time variation. Note that, outlier executions are important to study and while we focus on the common case in this work, we discuss power analysis for outlier executions in Section VI.
@@
Power profile differentiation: To tackle power variations (Figure 3d), we identify two specific power profiles of interest for a kernel based on our empirical experiments (Figure 4d). Specifically, we tag the power profile of the first kernel execution, post warm-up executions, beyond which the kernel execution time does not lower substantially (typically 3 warm-up executions from GPU idle state), as the steady-state execution (SSE) profile. This provides power visibility for the best execution time. Additionally, we tag the power profile of the kernel execution post SSE, beyond which the power does not vary substantially as the steady-state power (SSP) profile.
We observe that power can gradually rise from SSE to SSP and studying these edge points gives a potential range of power variation for a kernel in isolation, that is, when it is not perturbed by other kernels. That said, we don’t intend to convey that this range is strict or that the power of a kernel always falls in this range. As we discuss in Section V-C3, power manifested by a low-power kernel does depend heavily on kernels preceding the said kernel. As such, while we analyze both SSE/SSP profiles in this work, we also discuss the effect of interleaved kernel execution scenarios on SSE and SSP. Finally, note that, in some cases SSE and SSP profiles are the same execution/power profile. We discuss this further in Section V-D.
@@
IV-B FinGraV: Steps
  1. 1.
    Time the kernel a few times (4-5) to identify the kernel execution time. Use this to lookup the FinGraV empirical guidance table (Table I) to deduce the recommended #runs, binning margin, and a guidance on #LOIs to collect for the given kernel execution time.
  2. 2.
    Add relevant CPU-side instrumentation for GPU code. This includes timing the kernel start/end, reading the GPU timestamp before kernel execution (Section IV-A), and starting/ending power logger before/after the kernel.
  3. 3.
    For the SSE profile only, per run, execute the kernel four times. For the SSP profile, do a binary search starting at four executions per run (doubling or halving #executions) to ascertain the number of executions needed to attain the SSP profile. Note that an SSP run can also get the SSE profile.
  4. 4.
    Execute specified runs with specified executions per run.
  5. 5.
    Discard all but the golden runs. Golden runs are the ones that include SSE/SSP execution times belonging to the execution time bin with the maximum number of executions within binning margin of each other (for example 5% as shown in Table I for 25-200us entries).
  6. 6.
    Synchronize CPU-GPU time to identify LOI, if available, in the collected power logs per run. Identify also the TOI (the time in kernel execution where LOI was obtained).
  7. 7.
    If #LOIs obtained thus far are less than those suggested in Table I, optionally, execute more runs (= #LOIs).
  8. 8.
    Stitch the different runs by plotting all collected LOIs and TOIs.
@@ V FinGraV Profiles and Insights
We begin with a discussion of the AI computations under study in this work and our setup for executing them. We follow this by providing an evaluation of key tenets of FinGraV methodology and sharing some experimental profiling guidance. Finally, we discuss FinGraV profiles for AI computations, key observations from these profiles, and implications for future hardware/software based on these observations.
@@ V-A AI Operator Space and Setup
To execute GEMMs, we harness AMD ROCm [15] rocBLAS library [16] consisting of high-performance GEMM kernels. For AI collectives, we employ AMD ROCm Communication Collectives Library (RCCL) [17], a library of standard collective communication routines for GPUs. As discussed in Section IV-B, in a given run, we can execute a kernel multiple times and collect power logs over the entire run. We use post-processing to identify executions of interest within a given run.

### v2

# FinGraV: Methodology for Fine-Grain GPU Power Visibility and Insights

@@ Abstract
Ubiquity of AI makes optimizing GPU power a priority as large GPU-based clusters are often employed to train and serve AI models. An important first step in optimizing GPU power consumption is high-fidelity and fine-grain power measurement of key AI computations on GPUs. To this end, we observe that as GPUs get more powerful, the resulting sub-millisecond to millisecond executions make fine-grain power analysis challenging. In this work, we first carefully identify the challenges in obtaining fine-grain GPU power profiles. To address these challenges, we devise FinGraV methodology where we employ execution time binning, careful CPU-GPU time synchronization, and power profile differentiation to collect fine-grain GPU power profiles across prominent AI computations and across a spectrum of scenarios. Using the said FinGraV power profiles, we provide both, guidance on accurate power measurement and, in-depth view of power consumption on state-of-the-art AMD Instinct MI300X. For the former, we highlight a methodology for power differentiation across executions. For the latter, we make several observations pertaining to GPU sub-component power consumption and GPU power proportionality across different scenarios. We believe that FinGraV unlocks both an accurate and a deeper view of power consumption of GPUs and opens up avenues for power optimization of these ubiquitous accelerators.
@@ I Introduction
A key enabler for the current AI wave has been the compute and memory horsepower availed by modern GPUs. As the demand for AI sky-rockets, hyperscalers continue to deploy large-scale GPU clusters to train and serve AI models [1, 2]. As such, optimizing GPU power and resultant energy consumption can go a long way in reducing energy consumption and costs of AI deployments. Further, GPU power optimization is not only beneficial for AI but also for other domains, as the use cases for which GPUs are being utilized continue to widen. As an example, even for the HPC domain, the majority (more than 75%) of node-level power provisioning for Hewlett Packard Enterprise Frontier [3], the world’s first exascale supercomputer, is for GPUs.
An important first step in optimizing GPU power is better visibility into GPU power. Specifically, in this work, we focus on measuring fine-grain GPU power profiles along both time and space dimensions. That is, first, for fine-grain profiling along time, we focus on power measurements throughout the execution of a GPU computation (commonly referred to as a kernel in GPU parlance). Second, for fine-grain profiling along space dimension, we focus on breaking down GPU power consumption into sub-components (e.g., compute cores, memory, etc.).
Note that, we focus on accurate kernel-level power profiles for several reasons. First, as our work and prior work [4] show, large-scale ML compute kernels can often be GPU power limited. That is, hitting GPU power limit during kernel execution causes frequency throttling [5] leading to lower performance. Further, as applications are comprised of sequences of kernels, power-limited kernels essentially lead to lower performance at the application-level. Finally, as energy is simply power integrated over time, accurate fine-grain power profiles lead to accurate energy measurements at the application-level aiding in energy optimization as well. Overall, we believe that fine-grain power visibility is critical in identifying avenues for, and hence design of, intelligent GPU power management capabilities.
@@
However, fine-grain GPU power analysis is challenging for many reasons as depicted in Figure 1. First, as GPUs get more powerful (higher compute/memory throughputs), execution times for computations can often end up in the sub-millisecond (ms) to few ms range. This makes getting fine-grain power profiles challenging as natively available power samplers on GPUs often sample power at tens of milliseconds [6] (C1). Further, a common technique of repeated kernel executions is challenging to use as kernels with such low execution times can manifest execution time variation (e.g., due to slight differences in memory allocation and hence access patterns) which must be tackled (C2). Not only do execution time variations need to be accounted for, even power variations can occur across executions, across interleaving of various kinds of computations and more (C3) depending on averaging employed by underlying power samplers. Finally, using a high frequency GPU-side power sampler can either lead to repeated CPU-GPU communication for online measurements (as CPU schedules computations on the GPU today) or require careful synchronization of CPU-GPU time during post-processing of power logs obtained to identify logs that belong to the kernel execution (C4).
To tackle these challenges and provide fine-grain GPU power visibility, in this work, we propose FinGraV(Fine-Grain Visibility) methodology to collect fine-grain power profiles on the state-of-the-art AMD Instinct MI300X GPU, which is being used exclusively to serve all Llama 405B live traffic at Meta [7]. To design FinGraV, we harness GPU-side power logging and via careful synchronization of CPU-GPU time, identify power logs of interest to stitch together a fine-grain power profile in time. Further, we employ execution time binning strategy to tackle execution time variation. Finally, we provide guidance on tackling power variation that manifests across kernel executions by aiding the user in identifying the power profile of interest for a given kernel. Note that, as our work shows, it is critical for researchers to be aware of the said power variations as it can result in inferring inaccurate power/energy measurements (as high as 80% energy measurement error).
We study the profiles that FinGraV leads to for prominent AI computations (matrix-matrix multiplication and collective communication kernels) across a spectrum of scenarios. Using a spectrum of these profiles, we make several observations pertaining to GPU sub-component power consumption and GPU power proportionality via comparative analysis of different AI computations. We conclude with a discussion of power optimization opportunities that FinGraV profiles help identify.
@@
  • To address the above identified challenges, we propose FinGraV (Fine-Grain Visibility) methodology, where we employ execution time binning, careful CPU-GPU time synchronization, and power profile differentiation to collect fine-grain and accurate GPU power profiles.
@@
  • From the above FinGraV profiles, we make several observations pertaining to GPU sub-component power consumption and GPU power proportionality via comparative analysis of different AI computations.
@@ II-A AMD Instinct MI300X GPU and Sub-components
In this work, we focus on fine-grain power analysis of the state-of-the-art AMD Instinct MI300X GPU, based on AMD CDNA 3 architecture, shown in Figure 2. As depicted, the MI300X GPU is a chiplet-based design which harnesses advanced packaging to integrate heterogeneous chiplets, each specialized for a specific function. Reading the figure from the top, MI300X has accelerator complex dies (XCD) vertically stacked over I/O dies (IOD), which in turn are stacked over a passive silicon interposer. There are a total of four IODs each having two XCDs stacked for a total of eight XCDs in a single MI300X GPU [8, 9].
Figure 3: Challenges in doing fine-grain GPU power analysis.
@@ II-B ML Focus and Operators of Interest
While AI workloads comprise a variety of operators, we consciously focus on two primary operators for our analysis, namely, general matrix-matrix multiplications (GEMM) kernels and communication kernels for they contribute to the majority of the AI execution time [13]. We consider both GEMM and communication kernels across a spectrum of scenarios (GEMMs: compute versus memory-bound; communication: latency versus bandwidth-bound) in order to study their effect on power consumption.
@@ III Challenges to Fine-grain GPU Power Analysis
Low sampling frequency: As compute throughput and memory bandwidth made available by high-end GPUs continue to scale, short of problem sizes scaling commensurately, GPU kernel execution times can often end up in the sub-millisecond (ms) to few ms range. This is certainly true for the AI kernels we have benchmarked on MI300X GPU (Section V). This means that any GPU power measurements at low sampling frequency, as Figure 3a depicts, can completely miss sampling power for a given kernel.
@@
Power variance across executions: Finally, in addition to above execution time variation, underlying power loggers can employ averaging wherein multiple instantaneous power samples are averaged and reported at some time granularity. This in turn leads to power variations manifesting as well, which, if not tackled, can lead to considerable error in power and hence energy measurements. Specifically, as depicted in Figure 3d, we observe in this work that repeated executions of the same kernel (tagged 1/2/3 in Figure 3d) or interleaving of a kernel with other kernels (kernel 1 interleaved with kernel x and y) can lead to different power profiles. As such, identifying which power profiles to focus on is important.
@@ IV-A Addressing Challenges
In this section, we discuss the broad strokes of our solution to address the challenges we identified in Section III.
@@
On-GPU power logger: We harness a 1ms power logger available internally at AMD on MI300X; each power sample is the average of multiple instantaneous power readings in the last 1ms. We discuss using the proposed FinGraV methodology in tandem with externally available power logging tools on AMD GPUs such as amd-smi [14] in Section VI. Note that, due to the averaging nature of the available power logger, the profiles in this work provide a fine-grain time view of average power manifested at different points in the kernel execution.
@@
High-resolution CPU-GPU sync: As the internal power logger logs power measurements on the GPU while being agnostic of kernel start/end events as discussed in Section III, we need to employ careful syncing of CPU-GPU time to identify the power log which was taken during the execution of the kernel (LOI) and where in the kernel execution was the log taken (TOI). To do so, first, we read a GPU timestamp counter before the kernel execution from the CPU side. Second, we separately benchmark the delay to read this timestamp. Finally, along with power logging, we also log this timestamp value (depicted in Figure 4b). By associating the GPU timestamp with every power log and syncing a GPU timestamp (T0) with a specific timestamp on the CPU (Tc), we can post-process the power logs to identify LOI/TOI by identifying the kernel-start/end times (also in CPU time domain) in relation to the synced CPU time Tc.
@@
Kernel execution time binning: Note that, as discussed in Section III, a single run is insufficient to create fine-grain power profiles. That said, correlating measurements across runs is challenging due to the kernel execution time variation. To tackle this, we employ a simple strategy of execution time binning as depicted in Figure 4c. That is, based on our empirical experiments, we guide the user towards two heuristics for fine-grain power profiles: (1) #runs to execute and (2) margin of execution time variation to allow. By excluding outlier runs with considerable variations and fine-tuning the margin for kernel execution time (Section V-B), we lower the effects of execution time variation. Note that, outlier executions are important to study and while we focus on the common case in this work, we discuss power analysis for outlier executions in Section VI.
@@
Power profile differentiation: To tackle power variations (Figure 3d), we provide a methodology to differentiate power profiles in order to identify the most pertinent power profile for a kernel. That is, as the underlying power logger averages instantaneous power samples in the last 1ms, power gradually rises as we continually execute a given kernel. This happens because at the start, more idle power samples are averaged with the kernel’s power draw. However, with repeated executions, as the kernel executions fill up power averaging window of 1ms, a time series view of average power of the kernel emerges and power does not vary substantially beyond this point.
Consequently, we identify two specific power profiles as depicted in (Figure 4d). First, we tag the power profile of the first kernel execution, post warm-up executions, beyond which the kernel execution time does not lower substantially (typically three warm-up executions from GPU idle state), as the steady-state execution (SSE) profile. Without power profile differentiation as we suggest, this is the power profile that a typical user associates with a kernel. Additionally, we tag the power profile of the kernel execution post SSE, beyond which the power does not vary substantially as the steady-state power (SSP) profile. Note that, as power depends on voltage, frequency and temperature, SSP profile is by definition specific to a given voltage-frequency setting and as such can be affected by dynamic voltage-frequency scaling. That said, in this work, we did not control the decisions of the underlying power management infrastructure.
It is the SSP power profile that provides the time series view of average power at different points in kernel execution. Comparing SSP and SSE profiles, demonstrates the potential error in power (and hence energy) measurement that can manifest without the power profile differentiation that we recommend. Finally, note that, depending on how large the kernel execution time is in comparison to the power averaging time window (1ms for our setup), SSP and SSE profile can be the same.
IV-B FinGraV: Steps
  1. 1.
    Time the kernel five times to identify the kernel execution time. Use this to lookup the FinGraV empirical guidance table (Table I) to deduce the recommended #runs, binning margin, and a guidance on #LOIs to collect for the given kernel execution time.
  2. 2.
    Add relevant CPU-side instrumentation to GPU code. This includes timing the kernel start/end, reading the GPU timestamp before kernel execution (Section IV-A), and starting/ending power logger before/after the kernel.
  3. 3.
    For the SSE profile only, per run, execute the kernel four times. Note that, across all AI computations we studied, three executions sufficed for execution time stabilization. That said, this can be simply deduced for a given GPU/library/binary empirically.
  4. 4.
    For the SSP profile, ascertain the number of executions needed using: max (ceil [(averaging interval)/(kernel execution time)], executions for SSE), where the averaging interval for our setup is 1ms and executions for SSE are four. Note that an SSP run can also get the SSE profile. Note also that, should power (frequency) throttling incur during warmup runs (rise followed by fall of power), binary search can be necessary to deduce executions to get SSP profile.
  5. 5.
    Execute specified runs with specified executions per run. Note that, to attain power averages at unique TOI in kernel (Figure 3b), we add random delays before starting kernel executions per run.
  6. 6.
    Discard all but the golden runs. Golden runs are those that include SSP execution times belonging to the execution time bin with the maximum number of executions within binning margin of each other (for example 5% as shown in Table I for 25-200us entries).
  7. 7.
    Synchronize CPU-GPU time to identify LOI, if available, in the power logs collected per run. Identify also the TOI (the time in kernel execution where LOI was obtained).
  8. 8.
    If #LOIs obtained so far are less than those suggested in Table I, optionally, execute more runs (= #LOIs).
  9. 9.
    Stitch the different runs by plotting all collected LOIs and TOIs.
@@ V FinGraV Profiles and Insights
We begin with a discussion of AI computations under study in this work and our setup for executing them. We follow this by providing an evaluation of the key tenets of FinGraV methodology and sharing some experimental profiling guidance. Finally, we discuss FinGraV profiles for AI computations, key observations from these profiles, and implications for future hardware/software based on these observations.
@@ V-A AI Operator Space and Setup
To execute GEMMs, we harness AMD ROCm [15] rocBLAS library [16] (version 4.2.0) consisting of high-performance GEMM kernels. For AI collectives, we employ AMD ROCm Communication Collectives Library (RCCL) [17] (version 2.20.5), a library of standard collective communication routines for GPUs. As discussed in Section IV-B, in a given run, we execute a kernel multiple times and collect power logs over the entire run. We use post-processing to identify executions of interest within a run.